OpenStudy (anonymous):

find the slope of the line containing these points (1/4, -1/2) (3/4,1/2)

OpenStudy (amistre64):

(x, y) (3/4, 1/2) -(1/4,-1/2) ---------- (1/2, 1) slope = y/x = 1/(1/2) = 2